Write two unit multipliers f or the e quivalence: 60 min = 1 hour.
1
Expert's answer
2020-05-28T17:44:18-0400

1) 240 min = ((240 min) / 1) * ((1 hour) / (60 min)) = (240 * (1 hour) / (1 * 60)) = (240/60) hour = 4hour


2) 20 min = ((20 min) / 1) * ((1 hour) / (60 min)) = (20 * (1 hour) / (1 * 60)) = (20/60) hour = 1/3hour


If hours are converted into minutes a multiplier of 60 should be used. If minutes are converted into hours a multiplier of 1/60 should be applied.